A right cylinder has a fixed height of 6 units find the rate of change of the total surface area A with respect to r.

1
Expert's answer
2021-11-30T15:51:23-0500

surface area of cylinder:

"A=2\pi rh+2\pi r^2"

we have: "h=6"

then:

"A=12\pi r+2\pi r^2"


rate of change of the total surface area:

"dA/dr=12\pi+4\pi r"